OpenStudy (amistre64):

What would be the pros and cons of using a quotient rule integration by parts formula?

OpenStudy (amistre64):

\[\int\frac{t'}{b}=\frac{t}{b}+\int\frac{bt'}{b^2}\] or some such

myininaya (myininaya):

shouldn't it be \[\int\limits_{}^{}\frac{t'}{b}=\frac{t}{b}+\int\limits_{}^{}\frac{t}{b^2} ?\]

OpenStudy (amistre64):

maybe ... it was a quick write up :)

OpenStudy (amistre64):

\[\frac{t}{b}=\int\frac{t'}{b}-\int\frac{b't}{b^2}\]

myininaya (myininaya):

oh wait maybe i should have said b'

myininaya (myininaya):

ok i agree with your last equation

OpenStudy (amistre64):

:) good, since its just the quotient rule lol ive read that it is really just the product rule in a different form ..

OpenStudy (amistre64):

its workable, just havent seen it done alot if ever

myininaya (myininaya):

yes any quotient can be written as a product really no need for the quotient rule

OpenStudy (amistre64):

\[tb^{-1}=\int t'b^{-1}+\int tb'^{-1}\] \[tb^{-1}-\int tb'^{-1}=\int t'b^{-1}\]

myininaya (myininaya):

\[tb^{-1}=\int\limits_{}^{}t'b^{-1}-\int\limits_{}^{}b'b^{-2}t\] \[\int\limits_{}^{}t'b^{-1}=tb^{-1}+\int\limits_{}^{}b'b^{-2}t\]

OpenStudy (amistre64):

yeah, that looks contorted into shape :)

myininaya (myininaya):

i like product rule more

myininaya (myininaya):

i don't want to make one for quotient rule lol

myininaya (myininaya):

i like only \[(fg)'=f'g+fg'\] \[fg=\int\limits_{}^{}f'g+\int\limits_{}^{}fg'\] \[\int\limits_{}^{}fg'=fg-\int\limits_{}^{}f' g\]